a car travles a distance 45 mi with a speed of 60 how long did it take

Answer by Theo(13342)

rate * time = distance
distance = 45 miles
rate = 60 miles per hour.
formula becomes:
60 miles per hour * time = 45 miles.
divide both sides of this equation by 60 to get:
time = 45 miles / 60 miles per hour.
solve for time to get time = 45/60 hours = 3/4 of an hour.
convert to minutes to get time = 45 minutes.